Please Answer---a man moves on straight road for 8.4 km at 70 km/h then he walk 2 km for 30 min whats its avg velocity????
 

Grade:11

1 Answers

Ankit
7 years ago
Speed = Distance/Time

In case (1)
Speed = 70kmph
Distance = 8.4km
Time = Distance/Speed
Therefore, Time taken when he is running is = 8.4/70 = 0.12 hours
 
In case (2):
Distance Covered = 2km
Time taken = 30 minutes = 0.5 hours
 
Therefore, total number of hours taken = 0.5 + 0.12 = 0.62 hours
Total distance covered = 10.4km
 
Average Velocity = Total Displacement/Total Time Taken
 
Therefore, Average Velocity = 16.77kmph